# ASVspoof 5: Design, Collection and Validation of Resources for Spoofing, Deepfake, and Adversarial Attack Detection Using Crowdsourced Speech

### v2

@@ 1 Introduction
Born from a special session held at Interspeech in 2013, the efforts to develop detection solutions for speech data was initially spearheaded through the ASVspoof initiative and challenge series, founded over a decade ago [1, 2, 3, 4, 5, 6]. ASVspoof has evolved from a challenge in the detection of spoofing attacks implemented with now-legacy TTS and VC algorithms to encompass also the detection of deepfake attacks implemented with the most recent deepfake technology as well as their impact upon the reliability of ASV systems. The tracking of developments in the speech synthesis and voice conversion communities is key to the pursuit of reliable and robust detection solutions. Each challenge edition was hence accompanied by the collection and public release of an updated ASVspoof database. The most recent challenge edition, ASVspoof 5, is no exception. The design, collection and validation of such data is a non-trivial undertaking and requires coordinated and collaborative international effort; in this article we summarise the collective effort of 2.5 years of work conducted by members of 18 teams spanning three continents.
@@ 2.2 ASVspoof 5 database partitions
Figure 2: An illustration of the MLS English source database partitioning scheme and its use in generation of the ASVspoof 5 database, including training, development and evaluation sets illustrated in gray boxes. Numbers in the second row indicate the number of speaker in each partition. Those in rectangular blocks denote the number of utterances. The green box denotes the 8 TTS/VC systems used for generation of the ASVspoof 5 training set. The blue box denotes 8 TTS/VC systems used for generation of the development set whereas the red box denotes the 16 TTS/VC systems and adversarial attacks used for generation of the evaluation set. Solid black and coloured arrows denote use of data for the training of TTS/VC attacks or surrogate models. Dashed arrows denote the adaptation towards target speaker and attack generation using input text or source speech utterance.
@@ 2.4 Surrogate ASV, CM, and perceptual quality estimation systems
The surrogate systems are illustrated toward the lower middle of Figure 2 and are trained, developed and evaluated in the usual way. While the provision of surrogate systems [^11] To reduce the burden on data contributors, access to surrogate systems was provided through an online platform. Contributors could submit generated TTS/VC utterances and use estimated ASV/CM/perceptual quality scores for fine-tuning. offers an opportunity for attack algorithms to be fine-tuned or optimised, given that spoofed utterances are generated by independent contributors, it is not possible to estimate reliably the relative degree to which each attack is tuned, though some analysis is provided in Section 6. Invariably, tuning involves merely experimentation to verify the degree to which attacks are effective and to adapt the algorithms accordingly. Surrogate systems were used for fine-tuning attack algorithms that are later used in generation of ASVspoof 5 development and evaluation sets only (not for the train set).
@@ 2.5 Speaker and utterance selection
Data for all speakers in the evaluation set, including both targets and non-targets, are removed if data corresponding to the same speakers also appears in the Librispeech database [30]. Both Librispeech and MLS databases are sourced from LibriVox. Many popular self-supervised learning (SSL) models [31, Table IV] are trained using the Librispeech database and use of the same data within the ASVspoof 5 evaluation set would lead to biases in the evaluation of derived systems.
@@ 2.7 Challenge tracks
The ASVspoof 5 database was designed to support two different tasks which form a pair of challenge Tracks. As illustrated to the left of Figure 3, rack 1 is a spoof/deepfake detection task and concerns the design of CM systems which produce a score for each input utterance, wherein higher scores indicate a higher likelihood that the input utterance is bona fide (or, equivalently, a lower likelihood that it is spoofed). Track 1 does not involve the design of an ASV system. CMs are implemented using the ASVspoof 5 training (#2.1-#2.2) and development (#6.1-#6.3) sets with evaluation then being performed using the evaluation set (#7.1-#7.3).
@@ 3.1 Attack types
VC-based attacks are implemented using an input utterance containing the voice of a non-target speaker in place of input text, but again using one or more adaptation utterances which contain the voice of a target speaker. The voice in the input utterance is converted to that of the target speaker, whereas the linguistic content of the former is preserved. Approaches to VC can be categorized similarly to the TTS systems illustrated in Figure 4, though with input utterances in place of input text and with a waveform encoder in place of the text encoder. VC acoustic decoders typically use similar statistical models or DNNs as TTS, but with necessary modifications to accommodate input features of different dimension.
Adversarial attacks involve the introduction to an utterance of discreet perturbations designed specifically to increase CM/ASV error rates. Adversarial attacks usually take the form of subtle, additive noise, e.g., estimated via a fast gradient sign method [20], designed for a specific input utterance and targeted CM/ASV system [37, 38, 39]. These techniques can be costly in terms of computation, especially when adversarial noise is estimated for a large number of utterances. Though still designed for a specific CM/ASV system, less costly, more universal adversarial attacks, have also been reported [40, 41]. While these can be applied to any input, they are applied to already-spoofed utterances with the aim of making them more difficult to detect and are hence designed for the specific, underlying spoofing attack aker. Our use of such adversarial attacks is illustrated to the right of Figure 4, wherein post-processing adversarial modules are applied to spoofed utterances produced by a given TTS/VC system. Adversarial modules are trainable, convolutive filters which target a specific CM system [40] or a specific ASV system [41]. The trainable parameters (i.e., filter coefficients) are learned from the gradients back-propagated from the targeted CM/ASV system. The fine-tuning of a TTS/VC system according to the surrogate CM/ASV or perceptual quality estimator outputs (Section 2.4) is also a form of black-box adversarial training. In this case, the TTS/VC system observes scores or decisions rather than gradients.
@@ 3.2 Training set
A04: a neural TTS system based on Grad-TTS [43], featuring a diffusion-based acoustic decoder. The acoustic decoder generates a mel-spectrogram via an iterative reverse diffusion process [61]. From latent linguistic features generated by the text encoder, the acoustic decoder with a U-Net architecture [62] is used to predict a ‘residual’ (or, more specifically, the gradient field which maximizes the probability of data to be generated). The residual is then summed with the latent features, the output of which is again fed to the acoustic decoder to produce another residual. The process is repeated ten times, and the output of the last iteration is used as a generated mel-spectrogram. The text encoder is a Transformer encoder with text and speaker embedding inputs. Speaker embeddings are extracted using an ECAPA-TDNN speaker encoder. The vocoder is a HiFi-GAN.
@@
A08: a variational inference with adversarial learning for end-to-end text-to-Speech (VITS) [45] system which uses a Transformer-based text encoder and a HiFi-GAN vocoder. The acoustic decoder is based on normalizing flow, similar to A01-A03. However, the text encoder, the acoustic decoder, and the vocoder are jointly optimized by maximizing a variational evidence lower bound [45]. Speaker embeddings are x-vectors [63]. VITS produces especially high quality synthetic speech and is readily accessible with the open-sourced toolkit ESPNet [64].
@@ 3.4 Evaluation set
A17: a zero-shot TTS system, dubbed ZMM-TTS [51], originally designed for multi-speaker and multi-lingual applications. The text encoder comprises a BERT-like module [87] and a stack of feedforward blocks. The acoustic decoder, which uses a hierarchical vector-quantization VAE (VQ-VAE) [88] architecture [89], and the HiFi-GAN-based vocoder are jointly optimised. Training data is sourced from multiple datasets in six languages. [^14] The training data includes 2.4k utterances collected from 10 target and 2 non-target speakers, for whom there is also data in the ASVspoof 5 evaluation set but not overlapped with the 2.4k TTS training utterances. The use of public datasets across multiple application domains is now widespread meaning that data overlap is increasingly more difficult to avoid. Bearing in mind that this represents less than 2% of the speakers in the evaluation set, the overlap simulates a worse-case scenario whereby training is performed using data collected from the target speaker. Nevertheless, experimental validation (Section 7.2) shows that A17 is not more effective in attacking the ‘seen‘ target speakers than the ‘unseen‘. The speaker encoder is an ECAPA-TDNN pre-trained using the VoxCeleb 2 dataset [90].
@@
A24: a zero-shot VC system. A RNN-CNN waveform encoder is used to extract a phonetic posteriorgram [96, 82], a latent representation of mostly speaker-independent phonetic information. The acoustic decoder, also a CNN, combines the posteriorgam with ECAPA-TDNN speaker embeddings to predict target-speaker-dependent latent features. A HiFi-GAN vocoder is then used to produce speech in the voice of the target speaker. All components are trained in an end-to-end manner. A24 is similar to A16 in that input speech is disentangled into speaker-dependent and speaker-independent features. A key difference to A16 is use of GAN-based, end-to-end training.
A25: a zero-shot, diffusion-based VC system [53]. The Transformer-based waveform encoder extracts latent features resembling the mel-spectra of an ‘average voice’, i.e., mel-spectra averaged over the set of speakers in the training data. Latent features and speaker embeddings extracted from a RNN-GE2E speaker encoder [71] are then fed to the acoustic decoder and used as priors for a reverse diffusion process [61]. The resulting mel-spectrogram is then fed to a HiFi-GAN vocoder for waveform generation.
@@ 4.2 Shortcut artefacts
To illustrate the problem, consider a set of TTS systems which set the peak waveform amplitude to specific, pre-set levels. The peak amplitude is then an informative indicator of the content having been generated by one of the these same TTS systems. A detector which exploits the peak amplitude shortcut might perform well when it is assessed using similarly generated data, but will likely fail if, for example, a TTS system is configured to generate outputs with random peak amplitude. In reality, the peak amplitude is semantically unrelated to the detection problem; an utterance generated using TTS synthesis is a spoof/deepfake regardless of the peak amplitude.
@@
The post-processing pipeline is designed to reduce the differences between the artefact distributions for spoofed/deepfake and bona fide utterances. Let us assume an input waveform $\boldsymbol{x}=(x_{1},x_{2},\cdots,x_{N})$ with $N$ sampling points. First, the waveform is linearly scaled by a factor $r\in\mathbb{R}^{+}$ so that the peak amplitude $\max_{n}r\|{x}_{n}\|$ is equal to 1.0. Next, the duration of leading and trailing non-speech segments are trimmed, after annotation of segment boundaries using the Whisper system if the input is from the evaluation set, or an energy-based speech activity detector [105, §5.1] for the development set utterances. Given the annotated leading non-speech segment $(x_{1},x_{2},\cdots,x_{N_{s}})$ with $N_{s}$ samples, an index $\tilde{n}_{s}$ is drawn at random from the probability mass function $\text{Pr}(n)=\frac{\exp(-|x_{n}|)}{\sum_{i=1}^{N_{s}}{\exp(-|x_{i}|)}},n\in[1,N_{s}]$. The sub-segment $(x_{1},x_{2},\cdots,x_{\tilde{n}_{s}})$ is then trimmed. [^17] The time resolution of the segment boundary produced by the Whisper system (and the energy-based speech activity detector) is 10 ms on account of the stride of the input mel-spectrogram. The value of $N_{s}$ is hence a multiple of 160 (= 10 ms $\times$ 16 kHz). We draw $\tilde{n}_{s}$ from $\text{Pr}(n)$ so that there is randomness in the location of the trimming. $\text{Pr}(n)$ gives a higher chance to select the index at samples where the waveform amplitude is low. For the trailing non-speech segment, an index $\tilde{n}_{e}$ is drawn at random in the same way so that the sub-segment $(x_{\tilde{n}_{e}},x_{2},\cdots,x_{N})$ is trimmed. Finally, a float-valued duration is drawn at random from a uniform distribution between 4.0 and 10.0 seconds. A contiguous chunk of this duration is then selected from within the remaining speech segment as the pipeline output. There is no further scaling of the utterance energy since we observed similar energy distributions at the pipeline output for bona fide and spoofed/deepfake utterances without additional normalisation. All data is distributed in FLAC format with a sampling rate of 16 kHz. Immaterial metadata, such as the audiobook title, are removed lest it also serve as a shortcut to help distinguish between spoof/deepfake and bona fide utterances.
@@
Figure 6: Illustration of bona fide and spoofed samples from each attack in a t-SNE embedding space (sub-figure on the left). Dots in a grey colour correspond to the bona fide samples; others are spoofed samples. The contour encloses 98.9% of the probability mass of the bona fide samples estimated using a kernel density estimate method. The label of each data class is positioned at the mean of its corresponding data distribution. As alternative views, he top-right and bottom-right sub-figures, which plot all the spoofed samples in the same colour, show the data with and without codec/compression, respectively. The contours in solid lines encloses 98.9% bona fide samples in the sub-figures. For comparison, the contour from the sub-figure on the left side is plotted as dotted lines in the sub-figures on the right side.
@@ 5 Visualisation
We present a visualisation of the database characteristics using t-distributed stochastic neighbour embedding (t-SNE) plots [106] and a hierarchical clustering dendrogram [107, Section 14.3]. The objective is to highlight similarities and differences between bona fide and spoof/deepfake utterances.
Illustrated in Figure 6 is a representation of the ASVspoof 5 database in the form of a 2-dimensional t-SNE plot [106]. Each point corresponds to an embedding extracted using an ASV system (the baseline system described in § 7). Gray points correspond to bona fide utterances while coloured points correspond to one of the 32 attacks. To reduce computation time and visual clutter, points are shown for a random selection of 10% of the utterances in the training, development and evaluation sets, with and without encoding/compression for the evaluation set only. The recipe used to produce Figure 6 is the same as that used to produce similar plots in [108] and [8] and involves attack-level, within-class covariance normalisation, length normalisation and speaker-level data whitening. Also plotted as a gray solid line in Figure 6 is a confidence contour which encloses 98.9% of the probability mass of the bona fide data distribution. It is estimated using a kernel density estimate (KDE) method. [^18] We use the the Seaborn KDE API [109]. The confidence level of 98.9% was used in deriving a similar visualisation for the ASVspoof 2019 database [8, Figure 3], though in the form of a confidence ellipse.
Compared to a similar plot for the ASVspoof 2019 database presented in [8, Fig. 3], a greater proportion of the attacks lie within the confidence contour. However, attacks A07, A11, and A14 lie outside of the contour, suggesting distinct differences to bona fide utterances. A13, another outlier, collapses into a cluster with low variance, indicating a lack of inter-speaker variation. Illustrated to the right in Figure 6 is a pair of additional t-SNE plots, the difference between which shows the impact of encoding/compression. The contours are plotted in the same was as before. The overlap between bona fide and spoofed/deepfake utterances is greater with encoding/compression than without, indicating a greater challenge to detect spoofed/deepfake utterances.
@@
Figure 7 provides an alternative visualisation of the database in the form of a dendrogram obtained using an agglomerative hierarchical clustering approach, similar to that presented in [8, Figure 3] for the ASVspoof 2019 database. The dendrogram is derived from the same set of speaker embeddings used in generating the t-SNE visualisation. Implementation details are described in B. The clustering is based upon pairwise cosine similarities between embeddings extracted from bona fide utterances and spoofed/deepfake utterances generated using each of the 32 attacks, a visualisation of which is depicted in the colour-scaled heatmap also shown in Figure 7. The most similar utterances are indicated in dark blue, whereas the most different are indicated in dark red.

### v3

@@ 1 Introduction
Born from a special session held at Interspeech in 2013, the efforts to develop detection solutions for speech data were initially spearheaded through the ASVspoof initiative and challenge series, founded over a decade ago [1, 2, 3, 4, 5, 6, 7]. ASVspoof has evolved from a challenge in the detection of spoofing attacks implemented with now-legacy TTS and VC algorithms to encompass also the detection of deepfake attacks implemented with the most recent deepfake technology as well as their impact upon the reliability of ASV systems. The tracking of developments in the speech synthesis and voice conversion communities is key to the pursuit of reliable and robust detection solutions. Each challenge edition was hence accompanied by the collection and public release of an updated ASVspoof database. The most recent challenge edition, ASVspoof 5, is no exception. The design, collection and validation of such data is a non-trivial undertaking and requires coordinated and collaborative international effort; in this article we summarise the collective effort of 2.5 years of work conducted by members of 18 teams spanning three continents.
@@ 2.2 ASVspoof 5 database partitions
Figure 2: An illustration of the MLS English source database partitioning scheme and its use in generation of the ASVspoof 5 database, including training, development and evaluation sets illustrated in gray boxes. Numbers in the second row indicate the number of speakers in each partition. Those in rectangular blocks denote the number of utterances. The green box denotes the 8 TTS/VC systems used for generation of the ASVspoof 5 training set. The blue box denotes 8 TTS/VC systems used for generation of the development set whereas the red box denotes the 16 TTS/VC systems and adversarial attacks used for generation of the evaluation set. Solid black and coloured arrows denote use of data for the training of TTS/VC attacks or surrogate models. Dashed arrows denote the adaptation towards target speaker and attack generation using input text or source speech utterance.
@@ 2.4 Surrogate ASV, CM, and perceptual quality estimation systems
The surrogate systems are illustrated towards the lower middle of Figure 2 and are trained, developed and evaluated in the usual way. While the provision of surrogate systems [^11] To reduce the burden on data contributors, access to surrogate systems was provided through an online platform. Contributors could submit generated TTS/VC utterances and use estimated ASV/CM/perceptual quality scores for fine-tuning. offers an opportunity for attack algorithms to be fine-tuned or optimised, given that spoofed utterances are generated by independent contributors, it is not possible to estimate reliably the relative degree to which each attack is tuned, though some analysis is provided in Section 6. Invariably, tuning involves merely experimentation to verify the degree to which attacks are effective and to adapt the algorithms accordingly. Surrogate systems were used for fine-tuning attack algorithms that are later used in generation of ASVspoof 5 development and evaluation sets only (not for the train set).
@@ 2.5 Speaker and utterance selection
Data for all speakers in the evaluation set, including both targets and non-targets, are removed if data corresponding to the same speakers also appears in the Librispeech database [31]. Both Librispeech and MLS databases are sourced from LibriVox. Many popular self-supervised learning (SSL) models [32, Table IV] are trained using the Librispeech database and use of the same data within the ASVspoof 5 evaluation set would lead to biases in the evaluation of derived systems.
@@ 2.7 Challenge tracks
The ASVspoof 5 database was designed to support two different tasks which form a pair of challenge Tracks. As illustrated to the left of Figure 3, Track 1 is a spoof/deepfake detection task and concerns the design of CM systems which produce a score for each input utterance, wherein higher scores indicate a higher likelihood that the input utterance is bona fide (or, equivalently, a lower likelihood that it is spoofed). Track 1 does not involve the design of an ASV system. CMs are implemented using the ASVspoof 5 training (#2.1-#2.2) and development (#6.1-#6.3) sets with evaluation then being performed using the evaluation set (#7.1-#7.3).
@@ 3.1 Attack types
VC-based attacks are implemented using an input utterance containing the voice of a non-target speaker in place of input text, but again using one or more adaptation utterances which contain the voice of a target speaker. The voice in the input utterance is converted to that of the target speaker, whereas the linguistic content of the former is preserved. Approaches to VC can be categorized similarly to the TTS systems illustrated in Figure 4, though with input utterances in place of input text and with a waveform encoder in place of the text encoder. VC acoustic decoders typically use similar statistical models or deep neural networks (DNNs) as TTS, but with necessary modifications to accommodate input features of different dimension.
Adversarial attacks involve the introduction to an utterance of discreet perturbations designed specifically to increase CM/ASV error rates. Adversarial attacks usually take the form of subtle, additive noise, e.g., estimated via a fast gradient sign method [21], designed for a specific input utterance and targeted CM/ASV system [38, 39, 40]. These techniques can be costly in terms of computation, especially when adversarial noise is estimated for a large number of utterances. Though still designed for a specific CM/ASV system, less costly, more universal adversarial attacks, have also been reported [41, 42]. While these can be applied to any input, they are applied to already-spoofed utterances with the aim of making them more difficult to detect and are hence designed for the specific, underlying spoofing attack and/or speaker. Our use of such adversarial attacks is illustrated to the right of Figure 4, wherein post-processing adversarial modules are applied to spoofed utterances produced by a given TTS/VC system. Adversarial modules are trainable, convolutive filters which target a specific CM system [41] or a specific ASV system [42]. The trainable parameters (i.e., filter coefficients) are learned from the gradients back-propagated from the targeted CM/ASV system. The fine-tuning of a TTS/VC system according to the surrogate CM/ASV or perceptual quality estimator outputs (Section 2.4) is also a form of black-box adversarial training. In this case, the TTS/VC system observes scores or decisions rather than gradients.
@@ 3.2 Training set
A04: a neural TTS system based on Grad-TTS [44], featuring a diffusion-based acoustic decoder. The acoustic decoder generates a mel-spectrogram via an iterative reverse diffusion process [62]. From latent linguistic features generated by the text encoder, the acoustic decoder with a U-Net architecture [63] is used to predict a ‘residual’ (or, more specifically, the gradient field which maximizes the probability of data to be generated). The residual is then summed with the latent features, the output of which is again fed to the acoustic decoder to produce another residual. The process is repeated ten times, and the output of the last iteration is used as a generated mel-spectrogram. The text encoder is a Transformer encoder with text and speaker embedding inputs. The vocoder is a HiFi-GAN. The implementation is based on the code released by the original paper, but speaker embeddings are extracted using an ECAPA-TDNN speaker encoder.
@@
A08: a variational inference with adversarial learning for end-to-end text-to-Speech (VITS) [46] system which uses a Transformer-based text encoder and a HiFi-GAN vocoder. The acoustic decoder is based on normalizing flow, similar to A01-A03. However, the text encoder, the acoustic decoder, and the vocoder are jointly optimised by maximizing a variational evidence lower bound [46]. Speaker embeddings are x-vectors [64]. VITS produces especially high quality synthetic speech and is readily accessible with the open-sourced toolkit ESPNet [65].
@@ 3.4 Evaluation set
A17: a zero-shot TTS system, dubbed ZMM-TTS [52], originally designed for multi-speaker and multi-lingual applications. The text encoder comprises a BERT-like module [88] and a stack of feedforward blocks. The acoustic decoder, which uses a hierarchical vector-quantization VAE (VQ-VAE) [89] architecture [90], and the HiFi-GAN-based vocoder are jointly optimised. Training data is sourced from multiple datasets in six languages. [^14] The training data includes 2.4k utterances collected from 10 target and 2 non-target speakers, for whom there is also data in the ASVspoof 5 evaluation set but not overlapped with the 2.4k TTS training utterances. The use of public datasets across multiple application domains is now widespread meaning that data overlap is increasingly more difficult to avoid. Bearing in mind that this represents less than 2% of the speakers in the evaluation set, the overlap simulates a worse-case scenario whereby training is performed using data collected from the target speaker. Nevertheless, experimental validation (Section 7.2) shows that A17 is not more effective in attacking the ‘seen’ target speakers than the ‘unseen’. The speaker encoder is an ECAPA-TDNN pre-trained using the VoxCeleb 2 dataset [91].
@@
A24: a zero-shot VC system. An RNN-CNN waveform encoder is used to extract a phonetic posteriorgram [97, 83], a latent representation of mostly speaker-independent phonetic information. The acoustic decoder, also a CNN, combines the posteriorgam with ECAPA-TDNN speaker embeddings to predict target-speaker-dependent latent features. A HiFi-GAN vocoder is then used to produce speech in the voice of the target speaker. All components are trained in an end-to-end manner. A24 is similar to A16 in that input speech is disentangled into speaker-dependent and speaker-independent features. A key difference to A16 is use of GAN-based, end-to-end training.
A25: a zero-shot, diffusion-based VC system [54]. The Transformer-based waveform encoder extracts latent features resembling the mel-spectra of an ‘average voice’, i.e., mel-spectra averaged over the set of speakers in the training data. Latent features and speaker embeddings extracted from an RNN-GE2E speaker encoder [72] are then fed to the acoustic decoder and used as priors for a reverse diffusion process [62]. The resulting mel-spectrogram is then fed to a HiFi-GAN vocoder for waveform generation.
@@ 4.2 Shortcut artefacts
To illustrate the problem, consider a set of TTS systems which set the peak waveform amplitude to specific, pre-set levels. The peak amplitude is then an informative indicator of the content having been generated by one of these same TTS systems. A detector which exploits the peak amplitude shortcut might perform well when it is assessed using similarly generated data, but will likely fail if, for example, a TTS system is configured to generate outputs with random peak amplitude. In reality, the peak amplitude is semantically unrelated to the detection problem; an utterance generated using TTS synthesis is a spoof/deepfake regardless of the peak amplitude.
@@
The post-processing pipeline is designed to reduce the differences between the artefact distributions for spoofed/deepfake and bona fide utterances. Let us assume an input waveform $\boldsymbol{x}=(x_{1},x_{2},\cdots,x_{N})$ with $N$ sampling points. First, the waveform is linearly scaled by a factor $r\in\mathbb{R}^{+}$ so that the peak amplitude $\max_{n}r\|{x}_{n}\|$ is equal to 1.0. Next, the duration of leading and trailing non-speech segments are trimmed, after annotation of segment boundaries using the Whisper system if the input is from the evaluation set, or an energy-based speech activity detector [106, §5.1] for the development set utterances. Given the annotated leading non-speech segment $(x_{1},x_{2},\cdots,x_{N_{s}})$ with $N_{s}$ samples, an index $\tilde{n}_{s}$ is drawn at random from the probability mass function $\text{Pr}(n)=\frac{\exp(-|x_{n}|)}{\sum_{i=1}^{N_{s}}{\exp(-|x_{i}|)}},n\in[1,N_{s}]$. The sub-segment $(x_{1},x_{2},\cdots,x_{\tilde{n}_{s}})$ is then trimmed. [^17] The time resolution of the segment boundary produced by the Whisper system (and the energy-based speech activity detector) is 10 ms on account of the stride of the input mel-spectrogram. The value of $N_{s}$ is hence a multiple of 160 (= 10 ms $\times$ 16 kHz). We draw $\tilde{n}_{s}$ from $\text{Pr}(n)$ so that there is randomness in the location of the trimming. $\text{Pr}(n)$ gives a higher chance to select the index at samples where the waveform amplitude is low. For the trailing non-speech segment, an index $\tilde{n}_{e}$ is drawn at random in the same way so that the sub-segment $(x_{\tilde{n}_{e}},x_{2},\cdots,x_{N})$ is trimmed. Finally, a float-valued duration is drawn at random from a uniform distribution between 4.0 and 10.0 seconds. A contiguous chunk of this duration is then selected from within the remaining speech segment as the pipeline output. There is no further scaling of the utterance energy since we observed similar energy distributions at the pipeline output for bona fide and spoofed/deepfake utterances without additional normalisation. All data is distributed in FLAC format with a sampling rate of 16 kHz. Immaterial metadata, such as the audiobook title, is removed lest it also serve as a shortcut to help distinguish between spoof/deepfake and bona fide utterances.
@@
Figure 6: Illustration of bona fide and spoofed samples from each attack in a t-SNE embedding space (sub-figure on the left). Dots in a grey colour correspond to the bona fide samples; others are spoofed samples. The contour encloses 98.9% of the probability mass of the bona fide samples estimated using a kernel density estimate method. The label of each data class is positioned at the mean of its corresponding data distribution. As alternative views, the top-right and bottom-right sub-figures, which plot all the spoofed samples in the same colour, show the data with and without codec/compression, respectively. The contours in solid lines encloses 98.9% bona fide samples in the sub-figures. For comparison, the contour from the sub-figure on the left side is plotted as dotted lines in the sub-figures on the right side.
@@ 5 Visualisation
We present a visualisation of the database characteristics using t-distributed stochastic neighbour embedding (t-SNE) plots [107] and a hierarchical clustering dendrogram [108, Section 14.3]. The objective is to highlight similarities and differences between bona fide and spoof/deepfake utterances.
Illustrated in Figure 6 is a representation of the ASVspoof 5 database in the form of a 2-dimensional t-SNE plot [107]. Each point corresponds to an embedding extracted using an ASV system (the baseline system described in § 7). Gray points correspond to bona fide utterances while coloured points correspond to one of the 32 attacks. To reduce computation time and visual clutter, points are shown for a random selection of 10% of the utterances in the training, development and evaluation sets, with and without encoding/compression for the evaluation set only. The recipe used to produce Figure 6 is the same as that used to produce similar plots in [3] and [9] and involves attack-level, within-class covariance normalisation, length normalisation and speaker-level data whitening. Also plotted as a gray solid line in Figure 6 is a confidence contour which encloses 98.9% of the probability mass of the bona fide data distribution. It is estimated using a kernel density estimate (KDE) method. [^18] We use the Seaborn KDE API [109]. The confidence level of 98.9% was used in deriving a similar visualisation for the ASVspoof 2019 database [9, Figure 3], though in the form of a confidence ellipse.
Compared to a similar plot for the ASVspoof 2019 database presented in [9, Fig. 3], a greater proportion of the attacks lie within the confidence contour. However, attacks A07, A11, and A14 lie outside of the contour, suggesting distinct differences to bona fide utterances. A13, another outlier, collapses into a cluster with low variance, indicating a lack of inter-speaker variation. Illustrated to the right in Figure 6 is a pair of additional t-SNE plots, the difference between which shows the impact of encoding/compression. The contours are plotted in the same way as before. The overlap between bona fide and spoofed/deepfake utterances is greater with encoding/compression than without, indicating a greater challenge to detect spoofed/deepfake utterances.
@@
Figure 7 provides an alternative visualisation of the database in the form of a dendrogram obtained using an agglomerative hierarchical clustering approach, similar to that presented in [9, Figure 3] for the ASVspoof 2019 database. The dendrogram is derived from the same set of speaker embeddings used in generating the t-SNE visualisation. Implementation details are described in B. The clustering is based upon pairwise cosine similarities between embeddings extracted from bona fide utterances and spoofed/deepfake utterances generated using each of the 32 attacks, a visualisation of which is depicted in the colour-scaled heatmap also shown in Figure 7. The most similar utterances are indicated in dark blue, whereas the most different are indicated in dark red.